# Steering Language Model to Stable Speech Emotion Recognition via Contextual Perception and Chain of Thought

### v1

@@ 3.4 Implicit Chain-of-Thought
Implicit CoT Training. During this phase, we fine-tune the model on a combination of explicit and implicit CoT data. To ensure that the model maintains its ability to infer emotion categories from rich speech context during the self-distillation process, we gradually transition the training data from explicit to implicit CoT data using a linear schedule. Furthermore, we employ customized text prompts, illustrated in Figure 2 (stage 2), to guide the model reasoning process under the implicit framework. This approach enables C2SER to efficiently generate accurate emotion predictions while addressing the inefficiencies and error propagation associated with explicit CoT.
4 Experiment Setup
4.1 Datasets
Table 1: Speech corpora used to train C2SER.
[…]
ESD Zhou et al. (2021) Act Mix 5 31500 26.2
MER2024 Lian et al. (2024) TV Chinese 5 4414 5.9
MELD Poria et al. (2019) TV English 7 8851 7.4
BIIC-Podcast(V1.01) Upadhyay et al. (2023) Podcast Chinese 6 67186 141.3
MSP-Podcast(V1.12) Martinez-Lucas et al. (2020) Podcast English 6 156017 242.6
Internal dataset(Ours) / Mix 7 399369 785.3
Total - Mix 7 672668 1215.7
[…]
@@ 4.1 Datasets
The statistics of the training corpora are summarized in Table 1, which covers seven emotions: anger, happiness, neutral, sadness, surprise, disgust and fear. We utilize six open-source corpora that contain both emotion and text labels, including IEMOCAP Busso et al. (2008), MELD Poria et al. (2019), MSP-Podcast Martinez-Lucas et al. (2020), BIIC-Podcast Upadhyay et al. (2023), ESD Zhou et al. (2021), and MER2024 Lian et al. (2024), alongside an internal corpus containing text labels for model training. To obtain emotion labels for the internal corpus, we apply using an efficient automated labeling method using Emotion2Vec [^2] https://huggingface.co/Emotion2Vec/emotion2vec_plus_large for annotating speech emotions and GLM-4-9B-Chat [^3] https://huggingface.co/THUDM/glm-4-9b-chat Zeng et al. (2024) for annotating text emotions. We then take the intersection of the two annotations to ensure consistency and reliability. Additionally, we incorporate an internal speech corpus containing approximately 2400 hours of unlabeled data during the training of Emotion2Vec. The distributions of each emotion and the label construction for both explicit and implicit CoT are provided in Appendix 3 and Appendix B.
To comprehensively evaluate the model’s performance in downstream tasks, we follow the data allocation strategy of EmoBox Ma et al. (2024a) and use multiple publicly available datasets, including MELD Poria et al. (2019), ASVP-ESD Landry et al. (2020), EmoV-DB Adigwe et al. (2018), EMOVO Costantini et al. (2014), ESD, MESD Duville et al. (2021), M3ED Zhao et al. (2022), and CASIA Zhang and Jia (2008). Furthermore, considering the complexity of speech emotions in real-world scenarios, we introduce a diverse speech emotion test set. Specifically, we apply the automated labeling approach to annotate Emilia He et al. (2024), a large-scale multilingual and diverse speech generation resource with over 100,000 hours of speech data that captures a wide range of emotional contexts. We then manually verify the accuracy of the emotion labels. Each utterance is checked by at least two experts to ensure both accuracy and reliability. The final proposed test set, Emo-Emilia, consists of 1400 test samples, with 100 samples per emotion category across seven types in both Chinese and English (700 samples per language).
4.2 Implement Details
@@
The architecture of Emotion2Vec-S is based on the original Emotion2Vec [^4] https://huggingface.co/emotion2vec/emotion2vec_base model, with the addition of a classifier consisting of three fully connected layers. $\lambda_{utt}$ and $\lambda_{cate}$ are set to 0.1 and 100, respectively. We employ the Whisper-medium [^5] https://huggingface.co/openai/whisper-medium encoder for semantic feature extraction. The connection module is composed of a 4-layer Transformer followed by a linear layer, with intermediate feature dimensions set to 2,560 in the feed-forward module. For the text LLM component, we utilize the Qwen2-7B-Instruct [^6] https://huggingface.co/Qwen/Qwen2-7B-Instruct model Chu et al. (2024) and fine-tune it using Low-Rank Adaptation (LoRA) Hu et al. (2022). The LoRA rank is set to 8, the scaling factor is 32, and the dropout rate for LoRA matrices is 0.1.
@@
Table 2: Emotion2Vec-S performance on datasets of Chinese, English, Italian, Mexican and mixed languages. The best and the second best result is shown in bold and by underlined.
@@
4.3 Comparison Systems and Evaluation Metrics
We conduct comparative experiments with several advanced models, including self-supervised learning (SSL) models (e.g., WavLM Chen et al. (2022), Data2Vec Baevski et al. (2022), Data2Vec 2.0 Baevski et al. (2023)), and audio-language models (e.g., SECap Xu et al. (2024), Qwen2-Audio Chu et al. (2024), and SenseVoice-Small An et al. (2024)). For SSL models, we follow the methodology of EmoBox Ma et al. (2024a). First, features are extracted from the last Transformer layer of the pre-trained models and undergo uniform layer normalization to accelerate convergence. Then, a downstream network is applied to perform the SER task, which consists of a simple linear hidden layer, a ReLU activation function, a pooling layer, and a classification head. For ALMs, we train SECap on our speech corpora while directly using official checkpoints of Qwen2-Audio and SenseVoice-Small.
5 Experimental Results
5.1 Evaluation of Emotion2Vec-S
The results are presented in Table 2, where we compare Emotion2Vec-S with various SSL pre-trained models of similar model size and training corpora. Notably, Emotion2Vec-F refers to the Emotion2Vec model trained directly on the same corpora as Emotion2Vec-S, allowing us to investigate the impact of different datasets on model performance. The results demonstrate that Emotion2Vec-S consistently outperforms other models across most datasets. Interestingly, the performance gap between Emotion2Vec and Emotion2Vec-S indicates that the training corpus does influence the results, but it does not always lead to significant improvements. Nonetheless, Emotion2Vec-S consistently shows steady improvement compared to both Emotion2Vec-F and Emotion2Vec, validating the effectiveness of semi-supervised contrastive learning.
@@ 5.1 Evaluation of Emotion2Vec-S
Table 3: C2SER performance on datasets of Chinese, English, Italian, Mexican and mixed languages. Models are tested in a zero-shot fashion. The best and the second best result is shown in bold and by underlined.
@@
Qwen2-Audio’s results on EmoV-DB may indicate data leakage (i.e., inclusion of the dataset in training).
5.2 Evaluation of C2SER
We compare C2SER with several leading audio-language models (ALMs) across various test sets. As shown in Table 3, C2SER outperforms most comparison models on the majority of datasets. Notably, even with only 3,500 randomly preserved utterances for evaluation in the ESD dataset, C2SER still achieves exceptional performance. We speculate that this can be attributed to the shared origin of the training and test sets, a factor that may also explain the Qwen2-Audio results on EmoV-DB. However, when compared to SECap, which is trained on the same speech corpora, C2SER consistently outperforms SECap across all test datasets except MESD, where the performance of both models is similar. These results validate that incorporating speech context through the chain-of-thought approach effectively improves SER. Additionally, we observe significant improvements when comparing C2SER with explicit CoT to C2SER with implicit CoT. This performance gain highlights the success of self-distillation, which helps preserve reasoning capabilities while significantly reducing potential error accumulation in longer thought chains.
ALMs, including C2SER, also exhibit notable domain shifts across different languages. Specifically, C2SER outperforms comparison models on the M3ED and CASIA datasets, while achieving comparable results on MELD. Moreover, C2SER excels across all multilingual test sets. When extended to other languages, such as Italian, C2SER continues to lead. These findings demonstrate the strong language generalization capabilities of C2SER, indirectly suggesting that both contextual perception and chain-of-thought contribute to improving emotion recognition across diverse speech data.
5.3 Ablation Study
We conduct an ablation study to evaluate the contribution of each component in C2SER. The experimental results are presented in Table 4. Firstly, removing the Whisper encoder leads to a significant degradation in performance, with the model failing to converge during explicit CoT training due to the lack of semantic perception. Secondly, the model incorporating Emotion2Vec-S outperforms the version without it, demonstrating that acoustic perception is crucial for capturing emotional expressions effectively. Finally, excluding CoT causes a substantial drop in performance. This result suggests the reasoning capability of C2SER is improved after CoT training, which leads to a more accurate and stable emotion recognition.
Table 4: Ablation study of C2SER on the Emo-Emilia test set.
[…]
w/o Whisper encoder 32.07 32.07 34.56
w/o Emotion2Vec-S 57.93 57.93 51.10
w/o CoT 43.14 43.14 36.15
[…]
6 Conclusion
Limitation
Although C2SER effectively addresses the hallucination problem in ALM-based SER, we identified the need for further improvements in the performance across different languages. This is mainly due to the imbalance in the training corpora of C2SER. We believe that scaling up data size and leveraging balanced corpora to train C2SER will effectively improve performance across different languages.
Appendix A Data Distribution
The emotion and language distributions of the training corpora are shown in Figures 3 and 4. As observed, neutral emotions account for nearly half of the dataset, while fear and disgust constitute less than 2%. The scarcity of fear and disgust data arises from challenges such as subjective annotation (e.g., difficulty in accurate identification), limited natural occurrences in contextual expressions, and technical barriers in detecting these emotions from speech or text. Additionally, the proportion of Chinese speech is approximately double that of English speech. We plan to develop more balanced training corpora to address these disparities.
Figure 3: Distribution of each emotions.
Figure 4: Distribution of different language.
Appendix B Explicit CoT Construction
Table 5: Template for Explicit CoT data construction, label will be replaced with the corresponding value.
Prompt: Based on the provided speech features—including a speaking rate of <speaking rate label>, a volume level of <energy label>, and a pitch of <pitch label>—along with the text content ’<text label>’ and the emotion <emotion label>, generate a natural and logical emotional description. Here is an example: ’The speaker spoke at a <speaking rate label> pace, with a <pitch label> tone and <energy label> level: “<text label>”. Based on the analysis of speech characteristics, the emotion was inferred to be <emotion label>.’ Ensure including all speech features and logic of the description." Generated Example: The speaker spoke at a moderate pace, with a low-pitched tone and a soft volume: "together you sort of get this whole narrative of feedback where students are constantly feeling like you’re watching and you’re paying attention and you’re giving them suggestions I often think about like a basketball." Based on the speech characteristics, the emotion was inferred to be disgust, revealing a sense of resentment or aversion towards the described situation.
To construct training data for both explicit and implicit CoT, we first use signal processing tools to collect relevant speech attributes, followed by the generation of reasoning paths via a text-based LLM. Specifically, we compute pitch contours from speech using the PENN library  [^7] https://github.com/interactiveaudiolab/penn, and calculate the utterance-level mean. We extract energy from the speech using the pyloudnorm library  [^8] https://github.com/csteinmetz1/pyloudnorm, which returns a value representing the loudness level. The speaking rate is determined by dividing the number of phonemes in the transcript by the utterance length. Silences at both the beginning and end of the speech waveforms are trimmed for accuracy. After collecting statistics for all utterances, we calculate the mean $\mu$ and standard deviation $\sigma$, and divide the values into three levels based on the Central Limit Theorem: ’Low’ corresponds value below $\mu-\sigma$, ’Medium’ corresponds values between $\mu-\sigma$ and $\mu+\sigma$, and ’High’ corresponds values above $\mu+\sigma$.
After gathering speech attributes, we use GLM-4-9B-Chat to generate a CoT path conditioned on the speech attributes, original text transcripts, and emotion labels. A simple prompt is designed to guide GLM-4-9B-Chat to first produce speaking styles and speech content, followed by the final emotion. The template for this prompt is shown in Table 5.
Appendix C Evaluation Details
For all test sets, we retain utterances within the seven emotion categories: anger, happiness, neutral, sadness, surprise, disgust, and fear. Similar emotion labels, such as Amused, Joy, and Happy, are merged into the happiness category. During evaluation, SSL models are tested using leave-one-session-out five-fold cross-validation, with 20% of the training set used as the validation set for the Emo-Emilia test set. Evaluation on other test sets follows the EmoBox. Additionally, since the output of ALMs may include more than just emotion categories, we use Qwen2.5-14B-Chat [^9] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct to directly extract emotion labels from the descriptions generated by ALMs. The evaluation script is available in our repository.
To assess the performance of C2SER, we compare with the following systems.
Appendix D Category Accuracy of Emotion2Vec-S
Appendix E Category Accuracy of C2SER
Figure 6: Category Accuracy (%) of C2SER on the Emo-Emilia test set.
Appendix F Effectiveness of CoT training
Table 6: Effectiveness of CoT training on the Emo-Emilia test set.
[…]

### v2

@@ III-D Implicit Chain-of-Thought
Implicit CoT Training. During this phase, we fine-tune the model on a combination of explicit and implicit CoT data. To ensure that the model maintains its ability to infer emotion categories from rich speech context during the self-distillation process, we gradually transition the training data from explicit to implicit CoT data using a linear schedule. Specifically, we employ a batch-level mixing strategy where the probability of sampling an explicit CoT example decays linearly from 1.0 to 0.0 over the course of this training phase. This ensures that by the end of the training, the model is trained exclusively on implicit CoT data, fully internalizing the reasoning process. Furthermore, we employ customized text prompts, illustrated in Figure 2 (stage 2), to guide the model reasoning process under the implicit framework. This approach enables C2SER to efficiently generate accurate emotion predictions while addressing the inefficiencies and error propagation associated with explicit CoT.
IV Data Preparation
IV-A Training Data
TABLE II: Statistics of the preprocessed speech corpora used to train C2SER.
Dataset Source Emotion Labels Used Lang #Utts #Hrs
IEMOCAP Act Anger, Happiness, Neutral, Sadness English 5331 7.0
ESD Act Anger, Happiness, Neutral, Sadness, Surprise Mix 3500 29.1
MER2024 TV Anger, Happiness, Neutral, Sadness, Surprise Chinese 5030 5.9
BIIC-Podcast(V1.01) Podcast Anger, Happiness, Neutral, Sadness, Surprise, Disgust, Fear Chinese 70000 147.43
MSP-Podcast(V1.11) Podcast Anger, Happiness, Neutral, Sadness, Surprise, Disgust, Fear English 149307 237.94
Internal dataset(Ours) / Anger, Happiness, Neutral, Sadness, Surprise, Disgust, Fear Chinese 439300 788.35
Total - Anger, Happiness, Neutral, Sadness, Surprise, Disgust, Fear - 672668 1215.72
@@ IV-A Training Data
The statistics of the training corpora are summarized in Table II, which covers seven emotions: anger, happiness, neutral, sadness, surprise, disgust and fear. We utilize six open-source corpora that contain both emotion and text labels, including IEMOCAP [IEMOCAP], MELD [meld], MSP-Podcast [msp], BIIC-Podcast [biic], ESD [ESD], and MER2024 [mer2024], alongside an internal corpus containing text labels for model training. To obtain emotion labels for the internal corpus, we apply using an efficient automated labeling method using Emotion2Vec [^2] https://huggingface.co/Emotion2Vec/emotion2vec_plus_large for annotating speech emotions and GLM-4-9B-Chat [^3] https://huggingface.co/THUDM/glm-4-9b-chat [glm] for annotating text emotions. We then take the intersection of the two annotations to ensure consistency and reliability. Additionally, we incorporate an internal speech corpus containing approximately 2400 hours of unlabeled data during the training of Emotion2Vec. To ensure the quality and consistency of our training corpus, we applied a unified preprocessing pipeline to all datasets. This process included: (1) Label Filtering, where we kept only samples within our seven predefined emotion categories (e.g., merging ‘excited’ into ‘happiness’ for IEMOCAP); and (2) Duration Filtering, where we removed utterances longer than 20 seconds. Furthermore, for datasets with official splits like MELD, we used only the training set, and for ESD, we held out 3,500 samples as a separate test set.
The distributions of each emotion and the label construction for both explicit and implicit CoT are shown in Figures 4 and 5. As observed, neutral emotions account for nearly half of the dataset, while fear and disgust constitute less than 2%. The scarcity of fear and disgust data arises from challenges such as subjective annotation (e.g., difficulty in accurate identification), limited natural occurrences in contextual expressions, and technical barriers in detecting these emotions from speech or text. Additionally, the proportion of Chinese speech is approximately double that of English speech.
Figure 4: Training data emotion distribution: each slice represents a different emotion, with percentages shown.
Figure 5: Training data language distribution: each slice represents a different language, with percentages shown.
IV-B Emo-Emilia Dataset
Considering the complexity of speech emotions in real-world scenarios, we introduce a diverse speech emotion test set, Emo-Emilia. Specifically, we apply the automated labeling approach to annotate Emilia [emilia], a large-scale multilingual and diverse speech generation resource with over 100,000 hours of speech data that captures a wide range of emotional contexts. Specifically, we first perform automated annotation on speech and text modalities of the Emilia dataset: employing Emotion2Vec for speech emotion annotation and GLM-4-9B-Chat for text emotion labeling. We subsequently filtered samples with consistent emotion labels across both modalities to ensure annotation accuracy and reliability. From these aligned samples, we randomly selected 300 samples per emotion category covering both Chinese and English languages, resulting in 4,200 candidate data entries. Following this, four bilingual (Chinese-English) speech domain experts independently reviewed all samples, retaining only those with unanimous annotations from all reviewers to enhance data quality and consistency. The final test set, “Emo-Emilia”, comprises 1,400 samples, with 100 samples per emotion category across seven types in both Chinese and English (700 samples per language). The total duration of this test set amounts to approximately 3.3 hours, with 1.4 hours from Chinese speech and 1.9 hours from English speech. Detailed statistics are provided in Table III.
TABLE III: Statistics of the Emo-Emilia dataset. Each emotion category contains 100 Chinese (CN) and 100 English (EN) samples. Durations are in $H$:$M$:$S$ format.
Emotion # of samples CN Duration EN Duration
Anger 100CN+100EN 00:14:19 00:17:23
Happiness 100CN+100EN 00:12:19 00:17:52
Neutral 100CN+100EN 00:13:25 00:11:50
Sadness 100CN+100EN 00:13:14 00:20:08
Surprise 100CN+100EN 00:12:01 00:13:57
Disgust 100CN+100EN 00:10:35 00:15:46
Fear 100CN+100EN 00:14:19 00:14:38
Total 1400(700CN+700EN) 01:25:44 01:51:37
IV-C Evaluation Benchmarks
To comprehensively evaluate the model’s performance in downstream tasks, we follow the data allocation strategy of EmoBox [emobox] and use multiple publicly available datasets spanning multiple languages and usage scenarios: Chinese corpora CASIA [casia] and M3ED [m3ed], covering both studio-recorded and spontaneous speech; English corpora MELD [meld] and EmoV-DB [emov-db], representing conversational and acted settings; multilingual corpora ESD [ESD], ASVP-ESD [asvp-esd], and our proposed Emo-Emilia test set, which provide parallel Chinese–English annotations; and, for zero-shot cross-lingual evaluation, the Italian EMOVO [emovo] and Mexican MESD [mesd] datasets. This selection ensures broad linguistic diversity (monolingual to multilingual) and scenario diversity (acted, conversational, studio, spontaneous), enabling a systematic and robust assessment of model generalization across languages and real-world conditions. Detailed statistics for these evaluation datasets, including utterance counts, language distribution, and emotion categories, are summarized in Table IV.
TABLE IV: Details of the evaluation datasets.
Dataset Source Emotion Labels Lang #Utterances
CASIA Act Anger, Happiness, Neutral, Sadness, Surprise, Fear Mandarin 1200
M3ED TV Anger, Happiness, Neutral, Sadness, Surprise, Disgust, Fear Mandarin 24437
MELD TV Anger, Happiness, Neutral, Sadness, Surprise, Disgust, Fear English 13706
EmoV-DB Act Anger, Happiness, Neutral, Disgust, Sleepy English 6887
ESD Act Anger, Happiness, Neutral, Sadness, Surprise Mix 35000
ASVP-ESD Media Anger, Happiness, Neutral, Sadness, Surprise Mix 13964
EMOVO Act Anger, Happiness, Neutral, Sadness, Surprise, Disgust, Fear Italian 588
MESD Act Anger, Happiness, Neutral, Sadness, Disgust, Fear Mexican 862
Emo-Emilia Media Anger, Happiness, Neutral, Sadness, Surprise, Disgust, Fear Mix 1400
V Experiment Setup
V-A Implement Details
@@
The architecture of Emotion2Vec-S is based on the original Emotion2Vec [^4] https://huggingface.co/emotion2vec/emotion2vec_base model, with the addition of a classifier consisting of three fully connected layers. To ensure a fair comparison and isolate the gains from our proposed training method, we maintain the exact same backbone architecture, model size, and feature dimensions as the Emotion2Vec model. $\lambda_{utt}$ and $\lambda_{cate}$ are set to 0.1 and 100, respectively. The value of $\lambda_{cate}$ was empirically determined to balance the differing numerical scales of the loss components and was found to yield the best trade-off between training stability and discriminative performance. We employ the Whisper-medium [^5] https://huggingface.co/openai/whisper-medium encoder for semantic feature extraction. The connection module is composed of a 4-layer Transformer followed by a linear layer, with intermediate feature dimensions set to 2,560 in the feed-forward module. For the text LLM component, we utilize the Qwen2-7B-Instruct [^6] https://huggingface.co/Qwen/Qwen2-7B-Instruct model [Qwen2-Audio] and fine-tune it using Low-Rank Adaptation (LoRA) [lora]. The LoRA rank is set to 8, the scaling factor is 32, and the dropout rate for LoRA matrices is 0.1.
@@
V-B Comparison Systems and Evaluation Details
V-B1 Comparison Systems
We conduct comparative experiments with several advanced models. For self-supervised learning (SSL) models (e.g., WavLM [wavlm], Data2Vec [data2vec], Data2Vec 2.0 [data2vec2.0]), we follow the methodology of EmoBox [emobox]. First, features are extracted from the last Transformer layer of the pre-trained models and undergo uniform layer normalization to accelerate convergence. Then, a downstream network is applied to perform the SER task, which consists of a simple linear hidden layer, a ReLU activation function, a pooling layer, and a classification head. To ensure a fair comparison, we select models from EmoBox with a comparable parameter scale to that of Emotion2Vec-S, thereby controlling for model size in performance evaluation.
For ALMs, to assess the performance of C2SER, we compare with the following systems.
In addition to these end-to-end models, we introduce a strong cascaded system baseline to specifically validate the advantages of our unified framework.
Whisper-m + Qwen2-7B-Instruct: This system first transcribes speech into text using the Whisper-medium model and then feeds the plain text into a standalone Qwen2-7B-Instruct model for emotion recognition. This baseline represents an approach that relies solely on the semantic content of speech, allowing us to quantify the benefits of direct acoustic feature integration.
Our evaluation strategy for these systems varies based on their original design to ensure a fair and rigorous comparison. For foundation models with built-in SER functionality like Qwen2-Audio and SenseVoice-Small, we follow standard practice by directly evaluating their official checkpoints. Conversely, for a non-SER architecture like SECap, we train it on our speech corpora to establish a strong SER baseline.
V-B2 Evaluation Details
For all test sets, we first harmonize emotion labels into seven core categories (e.g., merging ‘amused,’ ‘joy,’ and ‘happy’ into the happiness category). The evaluation protocols then differ based on the model type: We test SSL models using leave-one-session-out five-fold cross-validation, following the EmoBox protocol. For ALMs, we conduct inference on designated test sets. Specifically, for datasets with official splits like MELD, we evaluate performance on the official test set. For the ESD dataset, we use the held-out set of 3,500 utterances that were not part of the training data. Since the output of ALMs can be free-form text, we then employ Qwen2.5-14B-Chat [^7] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct to directly extract the most appropriate emotion labels from the descriptions generated by ALMs. We use the following prompt: “Given the following text, determine its corresponding emotion and output only the single most appropriate emotion label. The possible labels are: anger, happiness, neutral, sadness, surprise, disgust, fear”
TABLE V: Emotion2Vec-S performance on datasets of Chinese, English, Italian, Mexican and mixed languages. The best and the second best result is shown in bold and by underlined.
@@ V-B2 Evaluation Details
VI Experimental Results
In this section, we first assess the quality of the Emotion2Vec-S representations by comparing them with other SSL pre-trained models. Next, we evaluate C2SER against leading audio-language models. We then examine category-level accuracies for both Emotion2Vec-S and C2SER, followed by an analysis of the impact of our chain-of-thought training. Finally, an ablation study dissects the contribution of each module within C2SER.
VI-A Evaluation of Emotion2Vec-S
The results are presented in Table V, where we compare Emotion2Vec-S with various SSL pre-trained models of similar model size and training corpora. Notably, Emotion2Vec-F refers to the Emotion2Vec model trained directly on the same corpora as Emotion2Vec-S, allowing us to investigate the impact of different datasets on model performance. The results demonstrate that Emotion2Vec-S consistently outperforms other models across most datasets. Interestingly, the performance gap between Emotion2Vec and Emotion2Vec-S indicates that the training corpus does influence the results, but it does not always lead to significant improvements. Nonetheless, Emotion2Vec-S consistently shows steady improvement compared to both Emotion2Vec-F and Emotion2Vec, validating the effectiveness of semi-supervised contrastive learning.
@@ VI-A Evaluation of Emotion2Vec-S
TABLE VI: C2SER performance on datasets of Chinese, English, Italian, Mexican and mixed languages. The best and the second best result is shown in bold and by underlined.
@@
Qwen2-Audio’s results on EmoV-DB may indicate data leakage (i.e., inclusion of the dataset in training). Results for MELD and ESD reflect in-domain evaluation. All other datasets were evaluated in a zero-shot, cross-dataset generalization setting.
VI-B Evaluation of C2SER
We compare C2SER with several leading audio-language models (ALMs) across various test sets, with the results presented in Table VI. C2SER consistently demonstrates superior performance over other end-to-end models like SECap, which was trained on the same corpora. This superiority is particularly evident when compared to the ‘Whisper-m + Qwen2-7B-Instruct’ cascaded baseline. The cascaded approach suffers from a severe performance degradation on datasets where acoustic features are dominant (e.g., ESD, CASIA), as it completely discards critical paralinguistic cues during the ASR step. In contrast, by processing acoustic and semantic information simultaneously, our end-to-end C2SER framework shows more robust and accurate emotion recognition across a wider range of scenarios. These results strongly validate that incorporating speech context through our unified, chain-of-thought approach effectively improves SER.
Furthermore, within our C2SER framework, we observe significant improvements when advancing from explicit to implicit CoT. This performance gain highlights the success of our self-distillation strategy, which preserves reasoning capabilities while significantly reducing the potential for error accumulation in longer thought chains.
Beyond comparing C2SER to other ALMs, it is crucial to position it relative to the traditional supervised paradigm. A comparison between Table VI and Table V reveals a noteworthy trend: on certain datasets (e.g., EmoV-DB and MESD), the zero-shot performance of C2SER is lower than that of Emotion2Vec-S evaluated with in-domain fine-tuning. This is not a model deficiency but an expected outcome stemming from their fundamental evaluation paradigms. Specifically, Emotion2Vec-S acts as a “domain expert” optimized for a specific dataset through supervised cross-validation. In contrast, C2SER acts as a “generalist model,” evaluated under zero-shot conditions on these datasets. Therefore, this comparison clearly illustrates the trade-off between performance and generalization: while a specialized “expert model” excels with sufficient in-domain data, the value of C2SER lies in its powerful zero-shot generalization capability, making it critically suited for real-world applications where data is sparse or domains constantly shift.
TABLE VII: Effectiveness of CoT training on the Emo-Emilia test set.
CoT Phase Model UA(%) ↑ WA(%) ↑ F1(%) ↑
Inference Qwen2-Audio 39.07 39.07 31.91